# Whose projection postulate?

### 23 Feb 2024 v1

@@ 4 Lüders
Gerhart Lüders misread von Neumann, claiming that von Neumann thought that the post-measurement state is always given by (3.1), with an orthonormal basis $\{|\psi_{\alpha i}\rangle\}$ determined by the measurement apparatus [^1] I was guilty of the same misreading in 1986, when I wrote “von Neumann stated a postulate about the results of measurement later than Dirac, and he got it wrong” [15].. Lüders validly argues that this would be absurd; as an extreme case, measurement of an observable represented by a multiple of the identity operator can be done without affecting the system at all, but what he attributes to von Neumann would be an extreme change of state, resulting in a totally mixed state, whatever the pre-measurement state. However, von Neumann’s caveat means that this post-measurement state applies only if the pre-measurement state is already totally mixed.
Lüders’s proposed amendment is precisely the projection postulate as stated in Section 1. It is now often called “the Lüders rule”, but as we have seen, it had already been stated by Dirac. Lüders claims that it “shows exactly what is meant by the expression ‘reduction of the wave function”’, for which the only reference he gives is Pauli’s Handbuch der Physik treatise [8]. We must assume that he was not familiar with Dirac’s book.

### 05 Mar 2024 v2

@@ 4 Lüders
Gerhart Lüders criticised von Neumann’s assumption that every measurement of a particular observable was associated with a complete orthonormal basis of each eigenspace of the observable, objecting that it led to “a most complicated mixture” after the measurement. This is certainly not always warranted; Lüders gave the extreme example that measurement of the unit operator can be achieved by doing nothing at all. There would then be no disturbance to the measured system, and there is no special orthonormal basis associated with the measurement.
Lüders proposed to replace von Neumann’s prescription by a single statement for each observable, irrespective of the method of measurement. This is the statement P3, which follows directly from the proposal of Dirac. Dirac acknowledged that this would not necessarily apply in every actual arrangement for measuring a given observable, but gave a reason (the “minimum disturbance”) for selecting this process for special consideration. Lüders, having noted this aspect of von Neumann’s discussion of measurement, gave it no further place in his own treatment.
The projection postulate, as stated in Section 1, is now often called “the Lüders rule”, but as we have seen, it had already been stated by Dirac. Lüders claims that it “shows exactly what is meant by the expression ‘reduction of the wave function”’, for which the only reference he gives is Pauli’s Handbuch der Physik treatise [8]. We must assume that he was not familiar with Dirac’s book.